NYSDEC Oversight

The NYSDEC provides oversight of activities conducted by the major responsible parties at various Greenpoint petroleum remediation sites in Brooklyn, New York. NYSDEC’s subcontractor (Ecology and Environment Engineering and Geology P.C. [E&E]) conducts monthly field oversight of PRP activities to assist NYSDEC in developing a better understanding of the day to day investigation and remedial activities that occur in the Greenpoint area and report any pertinent issues to NYSDEC’s project manager.

The following sections summarize recent activities observed at the various Greenpoint remediation sites in Brooklyn, New York.


Date of Oversight Observation: January 13, 2021

Former ExxonMobil Terminal

Description of Activities

  1. E&E observed the on-site treatment system in operation and recorded current operating parameters of the air strippers from the control screen.
  2. A contractor is working to install a 2,000-gallon settling tank for backwash water within the on-site treatment system building.
  3. E&E and Roux collected a sample of the on-site effluent for visual observation and compared it to a previously collected sample taken from the Newtown Creek in the vicinity of the outfall.

 Planned Activities

  1. Routine treatment system maintenance.

ExxonMobil Off-Site Plume Area

Description of Activities

  1. E&E personnel observed the off-site treatment system in operation.
  2. E&E observed that soil excavation activities along the 460 Kingsland Avenue bulkhead (OU3) have been completed.  
  3. E&E and Roux collected a sample of the off-site effluent for visual observation and compared it to a previously collected sample taken from Newtown Creek in the vicinity of the outfall.
  4. DMR sampling was completed Wednesday 1/13/21 prior to arrival onsite.

Planned Activities

  1. Routine treatment system maintenance.

KM Terminal (Former BP Terminal)

Description of Activities

  1. Antea conducted a site walkthrough with E&E personnel.
  2. Antea informed E&E that last week the OWS float switch for RW-6 was broken and RW-6 pumped LNAPL directly into the sump, which caused sheens to occur in Newtown Creek. The sheens were contained by booms. Daily sump cleanouts have been performed since detection and Antea stated that no sheens had been detected in Newtown Creek this week.
  3. Antea informed E&E that RW-5 and RW-6 recovery wells are off-line for turbidity testing today (1/13/21).
  4. Antea informed E&E that RW-11, RW-12, and RW-14 have measured increases in product. Antea believes this increase is due to interface meter issues. A new interface meter has been purchased and RW-11, RW-12, and RW-14 will be remeasured.
  5. E&E and Antea collected a sample of the on-site effluent for visual observation and compared it to a previously collected sample taken from the Newtown Creek in the vicinity of the outfall

 Planned Activities

  1. Routine treatment system maintenance.

Former Paragon Oil Terminal/Apollo Street Creek Parcels

Description of Activities

  1. E&E observed conditions along the upper, middle, and lower bulkheads, within the containment areas, and in Newtown Creek in the vicinity of the bulkhead. No sheens were observed in any of the containment areas or Newtown Creek.
  2. E&E personnel spoke with Arcadis via phone and was informed of the following:
    1. On 1/11/21 Arcadis began work inside the building during the annual winter shutdown.
    2. PID screening of all seams/cracks occurred within the building. The highest PID reading observed was 0.2 ppm.
    3. Sealing of small cracks/seams within the building began 1/11/21 and was completed 1/12/21; approximately 133 linear feet were repaired. 
    4. Removal of any LNAPL in wells which have more than 0.1 ft began on 1/12/21.
    5. The VPR system was repaired and restarted on 12/21/20.
    6. Quarterly sewer discharge sampling has been completed.
    7. System effluent discharge has been sampled for initial Newtown Creek discharge.
    8. Five drums of spent organic clay were taken offsite for disposal on 1/8/21. Spent carbon pick-up for offsite disposal of the waste is scheduled for 1/14/21.
    9.  Ambient air testing is scheduled for the week of 2/6/21, with 16-18 locations to be collected using 24-hour summa canisters.

 Planned Activities

  1. Spent carbon drum pickup for offsite disposal of the waste.
  2. Routine treatment system maintenance.

Newtown Creek Observations

Former Paragon Terminal Bulkhead
E&E observed conditions along the upper, middle, and lower bulkheads, within the containment areas, and in Newtown Creek in the vicinity of the bulkhead. No sheens or sediment plumes were observed.

Meeker Avenue Bulkhead
E&E observed the conditions of Newtown Creek at the end of Meeker Avenue in the vicinity of the ExxonMobil Off-site outfall and adjacent storm outfalls. Storm outfall D was observed not discharging water. No sheens or sediment plumes were observed in the vicinity or within visible distance upstream or downstream of the outfalls. This observation took place approximately 15 minutes after the 9:27 AM 4.9’ AMSL high tide.


This page was last updated on Monday, January 25, 2021.